# MapQaTor : A System for Efficient Annotation of Map Query Datasets

### v1

@@ 5 Conclusion
In this paper, we have proposed a novel framework, MapQaTor, first of its kind, to automatically fetch rich contextual map service data, which forms the basis to develop language-map benchmark datasets for evaluating SoTA LLMs. Our developed web platform simplifies data collection for users by offering precise spatial information, user-friendly search, and efficient data retrieval by using Map APIs. Our application also enables user to create geospatial questionnaire. Experimental evaluation suggests that MapQaTor is highly effective in developing geospatial question answer datasets. We believe this approach introduces a new task in geospatial question answering, which has the potential to open a new research direction in the intersection of language models and spatial reasoning. Furthermore, our framework can be adapted to other domains.
@@
Despite the capabilities of MapQaTor, several limitations should be acknowledged. The platform utilizes several paid map APIs, which may incur costs based on usage. During the current public demonstration period, users can explore its features without immediate expenses; however, in the long run, users will need to host the platform independently and integrate their own API keys to access paid functionalities. This requirement necessitates an understanding of the pricing structures associated with the various APIs, potentially impacting accessibility for some users. The platform’s functionality is heavily dependent on the availability and stability of external map APIs, meaning that any changes or deprecations in these APIs could negatively impact performance. The quality of the generated QA pairs is contingent on the retrieved data and users’ ability to formulate meaningful questions, which can introduce variability in dataset quality. The evaluation metrics used might not encompass all aspects of usability, possibly overlooking qualitative user feedback. In addition to map service data, other platforms such as Trip Advisor can also be a rich source of additional context for geospatial queries.
Appendix A Adapter Layer
The Adapter Layer in MapQaTor  provides base classes for various tools, including ‘TextSearch‘, ‘PlaceDetails‘, ‘NearbySearch‘, ‘ComputeRoutes‘, and ‘SearchAlongRoute‘. To integrate a new API under any tool, developers need to extend the specific base class and implement two methods: ‘convertRequest‘ and ‘convertResponse‘. These methods standardize the request and response formats to align with MapQaTor’s requirements, ensuring consistent interaction across APIs. Figure 8 demonstrates an implementation example of extending the ‘TextSearch‘ base class to support the TomTom API. Note that "key:TOMTOM_API_KEY" is used as a placeholder, which will be replaced in the backend using environment variables to prevent exposing the API key on the frontend.
@@ Appendix A Adapter Layer
Figure 8: Implementing the TomTom API Adapter for Text Search in MapQaTor
Appendix B Visualization Tools
For visualizing geospatial data, MapQaTor  utilizes the Google Maps JavaScript API to display places (see Figure 10) and routes (see Figure 11) directly on an embedded map. Markers are used in the map to represent places. It just needs the latitude and longitude of the place. On the other hand, routes are actually encoded polylines [^5] https://developers.google.com/maps/documentation/utilities/polylinealgorithm. But embeded map expects a set of coordinates along the path to visualize the route. So, we need to first decode the encoded polyline and convert it to a set of latitude and longitude. The decoding algorithm is shown in Figure 9.
function decodePolyline(polylineStr) {
let index = 0, lat = 0, lng = 0;
const coordinates = [];
const changes = { latitude: 0, longitude: 0 };
while (index < polylineStr.length) {
for (const unit of ["latitude", "longitude"]) {
let shift = 0;
let result = 0;
while (index < polylineStr.length) {
let byte = polylineStr.charCodeAt(index++) - 63;
result |= (byte & 0x1f)<<shift;
shift += 5;
if (byte < 0x20) break;
changes[unit] = result & 1 ? ~(result >> 1) : result >> 1;
lat += changes.latitude;
lng += changes.longitude;
coordinates.push({ lat: lat / 1e5, lng: lng / 1e5 });
return coordinates;
Figure 9: Polyline Decoding Algorithm
Figure 10: Set of markers indicating different places
Figure 11: Visualizing routes between places
Appendix C Exclusion of Temporal Variations in Routing APIs
To ensure reproducibility in the dataset, we have implemented specific measures to eliminate temporal variations in routing responses. These measures are essential for maintaining consistency across identical queries, regardless of when they are made.
Traffic Awareness Setting: In MapQaTor, routing APIs are set to "TRAFFIC_UNAWARE," meaning no real-time traffic data is used in generating routes and travel times. Real-time traffic data would introduce variability in travel times, as identical requests made at different times would yield different results based on current traffic conditions. By disabling traffic awareness, we ensure that each query consistently returns the same travel time, independent of real-time conditions, thereby enhancing dataset reproducibility.
Exclusion of Transit Mode: We have also excluded the "TRANSIT" travel mode from ’ComputeRoutes’ and ’SearchAlongRoutes’ tools. Transit routes, which include public transportation options like buses and trains, are heavily influenced by the specific time a request is made due to schedules, service frequency, and operational hours. Including transit mode would introduce further temporal variability, as the same query could yield different results depending on schedule changes, service frequency, and operational hours. This would undermine the consistency of the dataset, as identical queries could yield different travel times and routes.
Benefits of These Measures: By setting routes to be traffic-unaware and omitting transit mode, we ensure that:
  • Responses remain consistent: Identical queries yield the same routes and travel times, regardless of request timing.
  • Benchmarking is simplified: Model evaluations can focus purely on spatial and routing reasoning without accounting for real-time variations.
  • Reproducibility is maintained: The dataset provides a stable baseline for testing and comparing model performance over time.
These adjustments support our goal of creating a reproducible benchmark dataset, free from the temporal dependencies that could otherwise complicate consistent evaluation of geospatial reasoning tasks.
Appendix D Question Design and Annotation
D.1 Context Preview
During data collection process, users need to save the necessary fetched information, which we will refer to as Context. Context consists of a collection of API calls and their responses. After creating a context, annotators will proceed to create questions. They can view the created context in two mode.
  1. 1.
    Summarized Context: A textual summarized view as a list of fetched information. One liner for each information (see Figure 12).
    Figure 12: Summarized Context
  2. 2.
    Visual Context: In visual context, full details of each information in the context is shown with corresponding Map View ((see Figure 12).
    Figure 13: Visual Context
D.2 Place Name Suggestion
Annotators gather different places using TextSearch tool. At the time of writing question, on pressing ’@’ annotator will get the option to choose from available place names. On pressing enter, the selected name will be entered on the cursor position. This ensures the consistency of place names between context and questions in a hassle free way. This is also applicable in answer field.
Figure 14: Suggesting available places from the context
D.3 Answer Formats
We incorporate four distinct answer formats, which provide flexibility in how users interact with the system. These formats are crucial for creating a diverse set of training and evaluation data for large language models (LLMs). By supporting a range of response types, the dataset captures various ways that models must handle user queries and generate appropriate answers. The four answer formats are as follows:
  1. 1.
    Open Ended: The Open Ended format allows for unrestricted, free-text responses, enabling models to generate detailed and contextually rich answers. This format tests the model’s ability to understand and generate responses that may require creativity, explanation, or reasoning. It is particularly useful for questions that demand nuanced answers or elaboration.
    Figure 15: Answer format: Open Ended
  2. 2.
    Yes/No: The Yes/No format tests the model’s ability to provide binary, definitive answers. This format is valuable for questions that require a clear-cut response, such as confirming or denying information. The model’s performance in this format reflects its ability to handle factual questions with straightforward answers.
    Figure 16: Answer format: Yes/No
  3. 3.
    Multiple Choice: The Multiple Choice format allows the model to choose from a predefined set of possible answers, potentially selecting more than one option. This format challenges the model to identify the most relevant responses from a set of alternatives and is useful for assessing its ability to discern multiple correct answers or filter out incorrect options.
    Figure 17: Answer format: Multiple Choice
  4. 4.
    Single Choice: Similar to Multiple Choice, but with the restriction that only one option can be selected. This format is used to evaluate the model’s ability to determine the most appropriate single answer from a set of alternatives. It tests the model’s decision-making and accuracy in selecting the correct response.
    Figure 18: Answer format: Single Choice
D.4 Reference Field
After creating a QA pair, annotator need to select which of the informations in the context is actually relevant to this pair (see Figure 19). We are focusing on the informations fetched through PlaceDetails, NearbySearch, ComputeRoutes and SearchAlongRoute tool. This step is necessary to create a well-grounded and traceable dataset. This ensures that all the necessary information to answer a question is available in the context. Under the hood, corresponding api calls (Relevant API Calls) to each information are linked with the QA pair (see Fig. 35).
Figure 19: Choosing relevant informations to provide the correct answer.
Appendix E Example Annotation
On entering MapQaTor, we will see 3 major steps (see Fig. 20). First we need to design a context, which contains all the necessary information to design question-answer (QA) pairs. Then we will create QA pairs. And finally we can review and save the dataset.
Figure 20: Major steps of MapQaTor
As described in section 2.4, we can design context using 5 data collection tools (see Fig. 21).
Figure 21: Overview of data collection tools integrated into MapQaTor, showcasing essential functionalities.
Now let’s start designing a context. Let’s say we want to create questions on Louvre museum and Eiffel tower. For that we first need to search them using TextSearch tool and add them to context (see Fig. 22).
Figure 22: Example use of TextSearch tool
Next, we can search for the nearby restaurants of Eiffel Tower (see Fig. 23). Once we submit query, nearby places of Eiffel Tower will be added to the context and we can see the list of places with some key details (see Fig. 24).
Figure 23: Example use of NearbySearch tool
Figure 24: Nearby restaurants of Eiffel Tower
We can also search for available routes from Louvre Museum to Eiffel tower using ComputeRoutes tool (see Fig. 25). After submitting query, all the available routes with distance, duration and step-by-step navigation will be added to the context (see Fig. 26).
Figure 25: Example use of ComputeRoutes tool
Figure 26: Available routes from Louvre museum to Eiffel tower by car
Thus we can finish designing context, and move forward to create questions based on the context. We can start with an example question "Fastest time to go from Louvre Museum to Eiffel Tower by car?" and assign it to "Routing" category (see Fig. 27).
Figure 27: Example question creation (1)
Then we need to provide answer to the question. For our example question, we can get the answer from the context, which is "15 mins" (see Fig. 26). We can choose the answer type and provide the ground truth (see Fig. 28).
Figure 28: Example answer creation (1)
Upto this point, only the annotator knows which information in the context is used to derive the ground truth. To maintain traceability we need to link the QA pair with relevant information in the context. In this example, only the routing information from Louvre Museum to Eiffel Tower is relevant. Nearby Restaurants of Eiffel Tower is irrelevant. So, we will select "Driving route from Louvre Museum to Eiffel Tower" in "References" field (see Fig. 29).
Figure 29: Choosing relevant informations (1)
Similarly, we can add a second question "Nearby restaurant of Eiffel Tower with over 2000 user ratings and at least 4.5* rating" (see Fig. 30), provide answer (see Fig. 31) and link relevant informations (see Fig. 32).
Figure 30: Example question creation (2)
Figure 31: Example answer creation (2)
Figure 32: Choosing relevant informations (2)
Finally, we can review the context and QA pairs (see Fig. 33) and download the dataset in JSON format (see Fig. 34). Downloaded dataset format is shown in Figure 35.
Figure 33: Review context and QA pairs
Figure 34: Download dataset as JSON
"questions": [
"title": "Fastest time to go from Louvre Museum to Eiffel Tower by car?",
"answer": {
"type": "single-choice",
"options": [
"14 mins","15 mins",
"16 mins","17 mins"
"correct": 1
"classification": "routing",
"relevant_api_calls": [ 98006 ]
"title": "Nearby restaurant of Eiffel Tower with over 2000 user ratings and at least 4.5* rating.",
"answer": {
"type": "single-choice",
"options": [
"Cafe Fleur","Jules Verne",
"Linette","LAbreuvoir"
"correct": 3
"classification": "nearby",
"relevant_api_calls": [ 12492 ]
"api_call_logs": [
"url": "https://routes.googleapis.com/directions/v2:computeRoutes",
"method": "POST",
"headers": { .... },
"data": {
"origin": { .... },
"destination": { .... },
"travelMode": "DRIVE",
"uuid": 98006,
"result": { .... }
"url": "https://places.googleapis.com/v1/places:searchText",
"method": "POST",
"headers": { .... },
"data": {
"locationBias": { .... },
"includedType": "restaurant",
"uuid": 12492,
"result": { .... }
Figure 35: Example dataset where each QA pair is linked with relevant api calls

### v2

@@ 5 Conclusion
In this paper, we have proposed a novel framework, MapQaTor, first of its kind, to automatically fetch rich contextual map service data, which forms the basis to develop language-map benchmark datasets for evaluating SoTA LLMs. Our developed web platform simplifies data collection for users by offering precise spatial information, user-friendly search, and efficient data retrieval by using Map APIs. Our application also enables user to create geospatial questionnaire. Experimental evaluation suggests that MapQaTor is highly effective in developing geospatial question answer datasets. We believe this approach introduces a new task in geospatial question answering, which has the potential to open a new research direction in the intersection of language models and spatial reasoning.
@@
Despite the capabilities of MapQaTor, several limitations should be acknowledged. The platform utilizes several paid map APIs, which may incur costs based on usage. During the current public demonstration period, users can explore its features without immediate expenses; however, in the long run, users will need to host the platform independently and integrate their own API keys to access paid functionalities. This requirement necessitates an understanding of the pricing structures associated with the various APIs, potentially impacting accessibility for some users. The platform’s functionality is heavily dependent on the availability and stability of external map APIs, meaning that any changes, deprecations, or invalid API keys can negatively impact performance. The quality of the generated QA pairs is contingent on the retrieved data and users’ ability to formulate meaningful questions, which can introduce variability in dataset quality. The evaluation metrics used might not encompass all aspects of usability, possibly overlooking qualitative user feedback. In addition to map service data, other platforms such as Trip Advisor can also be a rich source of additional context for geospatial queries.
Appendix A Data Collection Tools
Figure 4: Search for a place
Figure 5: Fetch full details of a place
Figure 6: Search Nearby Places
Figure 7: Find routes between places
Figure 8: Search places along a route
Appendix B Prompt Design Interface
The prompt design interface enables users to generate prompts for LLM evaluation by selecting a structured or formatted context. It displays the generated prompt, ground truth answers, and Gemini’s response for comparison. Figure 9 illustrates this process.
Figure 9: The figure illustrates prompt creation, ground truth comparison, and Gemini’s response assessment.
Appendix C Exclusion of Temporal Variations in Routing APIs
To ensure reproducibility, MapQaTor  removes temporal variations in routing by:
Traffic Awareness Setting: Routing APIs are set to "TRAFFIC_UNAWARE," ensuring consistent travel times by ignoring real-time traffic.
Exclusion of Transit Mode: The "TRANSIT" mode is excluded to prevent variability from schedule changes.
Benefits:
  • Ensures consistent responses for identical queries.
  • Focuses evaluations on spatial reasoning, not real-time changes.
  • Provides a stable baseline for model benchmarking.
These measures enable reliable and reproducible geospatial evaluations in MapQaTor .
Appendix D API Extension Mechanism
Figure 11 demonstrates how new map services are integrated by extending MapQaTor ’s core tools:
Appendix E Place Name Suggestion
Using the TextSearch tool, annotators can retrieve place names. While writing a question or answer, pressing ’@’ suggests available place names, ensuring consistency between context and QA pairs.
Figure 10: Suggesting available places from the context.
@@ Appendix E Place Name Suggestion
Figure 11: Extending Text Search for TomTom API